OpenStudy (anonymous):

Solve for z: x+2y+x= 4 x+2z=3 -y-4z=-5

OpenStudy (***[isuru]***):

is the 1st one = x + 2y + x =4 or x + 2y +z = 4

OpenStudy (anonymous):

ah, so sorry the first one is x+2y+ Z

OpenStudy (***[isuru]***):

Hi, then let's number the equations first, x + 2y + z = 4 ---(1) x + 2z = 3 ---(2) -y - 4z = -5 ---(3) let's multiply (3) by -2 2y + 8z = 10 ---(4) (2) + (4) x + 2y + 10z =13 ---(5) (5) - (1) 9z = 9 z = 1 hope this will help u out